OpenStudy (anonymous):

Between which two consecutive number does the square root of 7 lie

OpenStudy (anonymous):

2^2=4,\[\sqrt{4}=2\] 3^2=9 \[\sqrt{9}=3\] \[so~\sqrt{7}~lies~\between~2~and~3\]